# Deflickering Vision-Based Occupancy Networks through Lightweight Spatio-Temporal Correlation

### v3

@@ IV-C Comparison Results
We compare our method with several state-of-the-art models, including single-frame VON methods such as Atlas [54], TPVFormer [37], MonoScene [18], and SurroundOcc [12], as well as history-aware VON methods such as BEVFormer [55] and BEVDet4D-Occ [13]. To ensure a fair comparison, all baselines and OccLinker are trained on the same ground truth data using identical training protocols. As shown in Table III, integrating OccLinker into existing baselines like MonoScene [18] and SurroundOcc [12] leads to notable performance gains. In particular, the combination of OccLinker with SurroundOcc achieves the best results among all models, improving IoU and mIoU (All) by 5.18% and 1.82%, respectively.
Occupancy accuracy on Occ3D
We also evaluate our method on the Occ3D benchmark [6], as shown in Table IV. The baselines include single-frame VON methods such as MonoScene [18], OccFormer [1], SurroundOcc [12], and ViewFormer [16], as well as history-aware VON methods including FB-Occ [17], SparseOcc [15], BEVDet4D-Occ [13], and OPUS-L [14]. When integrated into single-frame VONs, OccLinker consistently improves the performance of the base models. In particular, it boosts the performance of ViewFormer by 0.34% in IoU and 2.08% in mIoU. Furthermore, ViewFormer+OccLinker surpasses all existing history-aware VON methods [14, 56, 15, 17] by more than 2% in mIoU, demonstrating the effectiveness of OccLinker in modeling temporal information.
Method IoU $\uparrow$ mIoU $\uparrow$ $\overline{S_{m}}\uparrow$ $\overline{S_{s}}\uparrow$
@@
Atlas [54] 28.66 15.00 12.64 17.35
TPVFormer [37] 30.86 17.10 14.04 20.15
BEVFormer [55] 30.50 16.75 14.17 19.33
BEVDet4D-Occ [13] 24.26 14.22 11.10 17.34
MonoScene [18] 10.04 1.15 0.24 2.07 46.53 81.77
MonoScene [18] +OccLinker ($\Delta$) 13.10 +3.06 1.69 +0.54 0.34 +0.10 3.04 +0.98 54.21 +7.68 83.84 +2.07
SurroundOcc [12] 31.49 20.30 18.39 22.20 58.33 91.71
SurroundOcc [12] +OccLinker ($\Delta$) 33.12 +1.63 20.67 +0.37 18.26 -0.13 23.08 +0.88 60.64 +2.31 92.54 +0.83
Table III: Occupancy prediction performance on the SurroundOcc benchmark [12]. Performance improvements ($\Delta$) are highlighted in bold, and the best results are marked with .
Method IoU $\uparrow$ mIoU $\uparrow$ $\overline{S_{m}}\uparrow$ $\overline{S_{s}}\uparrow$
@@
MonoScene [18] 6.06 5.36 6.68
OccFormer [1] 21.93 21.78 22.06
FB-OCC [17] 39.11 33.74 43.88
SparseOcc [15] 30.10
BEVDet4D-Occ [13] 39.30 29.09 42.16
OPUS-L [14] 36.20 31.25 40.44
SurroundOcc [12] 51.89 7.24 0.36 13.35 65.35 89.54
SurroundOcc [12] +OccLinker ($\Delta$) 52.13 +0.24 10.3 +3.09 1.98 +1.62 17.76 +4.41 69.60 +4.25 90.91 +1.37
ViewFormer [16] 70.39 40.46 33.73 46.45 67.26 86.06
ViewFormer [16] +OccLinker ($\Delta$) 70.63 +0.24 41.30 +0.84 34.33 +0.60 47.50 +1.05 70.13 +2.87 87.10 +1.04
Table IV: Occupancy prediction performance on the Occ3D benchmark [12]. Performance improvements ($\Delta$) are highlighted in bold, and the best results are marked with .
@@
The $\overline{S_{m}}$ and $\overline{S_{s}}$ results presented in Tables III and IV show that integrating OccLinker consistently improves the temporal consistency of occupancy predictions across all frames and scenes for all models, highlighting the effectiveness of our approach. This improvement can be attributed to the use of previous static features and the incorporation of motion information, which together provide valuable correlations between static/dynamic cues and occupancy refinement.
@@ IV-D Ablation study
Table V presents the ablation results of different decomposed correlation designs in OccLinker, as illustrated in Fig. 4, when integrated with SurroundOcc [12] under $L=1$. M0 denotes the baseline model without OccLinker. M1 removes the correlation between historical and current static features. M2 omits the current static features entirely. M3 excludes the correlation between motion features and current static features. M4 reverses the roles of queries and key/value pairs in the correlation design of Fig. 4. M4 represents the complete version of OccLinker with all components included.
@@
Table VI investigates the impact of different types of motion information on SurroundOcc+OccLinker when $L=1$. Specifically, I0 denotes the base model [12] without any motion input. I1 uses raw high-frequency intermediate frames as motion input for OccLinker. I2 incorporates optical flow [49], while I3 employs frame difference [42] to extract motion cues.
Idx. Motion type IoU $\uparrow$ mIoU $\uparrow$ $\overline{S_{m}}\uparrow$ $\overline{S_{s}}\uparrow$ Time $\downarrow$
@@
As shown in Table VI, I1 improves IoU over I0, but performs worst in mIoU, $\overline{S_{m}}$, and $\overline{S_{s}}$ among the motion-aware variants. This performance drop is largely due to the presence of redundant and irrelevant information in raw intermediate frames, which complicates the extraction of motion features by OccLinker. Moreover, I3 significantly outperforms I2 in both IoU and mIoU, and achieves slightly better results in $\overline{S_{m}}$ and $\overline{S_{s}}$. This suggests that frame difference is more effective at capturing abrupt scene changes, such as pedestrians suddenly appearing or vehicles exiting intersections, whereas optical flow may introduce delays in responding to such dynamics. In addition, thanks to the lightweight design of OccLinker, using frame difference further reduces computational overhead by operating on simple differential signals, thereby enhancing inference efficiency.
@@
We conduct ablation experiments on the temporal context window length $L\in\{0,1,2,3,4\}$ in $\mathcal{W}$. The case of $L=0$ corresponds to SurroundOcc [12], which does not utilize any historical keyframes. As shown in Table VII, the best IoU and mIoU are achieved when $L=1$, while the highest $\overline{S_{m}}$ and $\overline{S_{s}}$ scores occur at $L=2$.
Idx. Window length $L$ IoU $\uparrow$ mIoU $\uparrow$ $\overline{S_{m}}\uparrow$ $\overline{S_{s}}\uparrow$
@@
Increasing $L$ enables the model to access more historical information, which can enrich temporal context. However, larger values of $L$ also introduce more irrelevant or noisy features, weakening the effectiveness of correlation modeling and potentially causing semantic misalignment. As a result, performance in IoU and mIoU begins to degrade when $L>2$ due to the accumulation of such interference.
@@ IV-E Case analysis
To qualitatively assess the effectiveness of our method, we compare SurroundOcc+OccLinker with the a single-frame 3D VON baseline (SurroundOcc [12]) and a history-aware VON method (BEVDet4D-Occ [13]).
@@
As shown in Fig. 5 (Scene 277, Frames #7–#11), a pedestrian walking along the sidewalk – parallel to the ego vehicle’s trajectory – is intermittently occluded by roadside vegetation. SurroundOcc [12] exhibits significant instability, with the pedestrian missing in Frames #7 and #9, highlighting its limitations in temporal modeling. BEVDet4D-Occ [13] partially alleviates this issue by fusing historical latent features into the current latent space, but still shows occasional inconsistencies, such as a detection dropout in Frame #8. In contrast, SurroundOcc+OccLinker completely eliminates these flickering artifacts and achieves consistent detection across all occlusion states.
@@
Fig. 7 showcases several challenging scenarios: (i) Vehicle-pedestrian occlusion (Scene-0911 Frame #15, Scene-0928 Frame #14): Both SurroundOcc [12] and BEVDet4D-Occ [13] fail to recover the occupancy state of the partially occluded pedestrian, whereas our method accurately localizes the target with precise geometry. (ii) Ambiguous road boundary (Scene-0923 Frame #28): Baseline methods produce fragmented or erroneous occupancy predictions. In contrast, our approach correctly anticipates the right-turn road geometry and preserves superior shape consistency with the actual road layout.
@@ IV-F Efficiency–performance tradeoff analysis
To ensure a fair comparison, all overhead analysis experiments are conducted on a single NVIDIA L20 GPU. Table Table IX summarizes the results of several state-of-the-art methods, including history-aware approaches (FB-Occ [17], SparseOcc [15], OPUS-L/M/S/T [14], and BEVDet4D-Occ [13]), the single-frame 3D VON baseline (ViewFormer [16]), and our method combined with ViewFormer. All models are evaluated on the Occ3D benchmark [6] using the same 2D visual encoder backbone, ResNet-50, which outputs feature maps with a resolution of $(32,88)$. The table reports the mIoU, GPU memory usage during inference, and per-sample latency of each method.
@@
FB-Occ [17] 39.11 5,933 MB 0.09 s
SparseOcc [15] 30.10 7,147 MB 0.05 s
OPUS-L [14] 36.20 10,579 MB 0.16 s
OPUS-M [14] 35.60 8,912 MB 0.11 s
OPUS-S [14] 34.20 7,320 MB 0.07 s
OPUS-T [14] 33.20 6,711 MB 0.03 s
BEVDet4D-Occ [13] 39.30 4,689 MB 0.26 s
ViewFormer [16] 40.46 4,582 MB 0.11 s
ViewFormer [16] +OccLinker ($\Delta$) 41.30 +0.84 4,687 MB +105 MB 0.12 s +0.01 s

### v4

@@ IV-C Comparison Results
We compare our method with several state-of-the-art models, including single-frame VON methods such as Atlas [Atlas], TPVFormer [TPVFormer], MonoScene [MonoScene], and SurroundOcc [surroundOcc], as well as history-aware VON methods such as BEVFormer [BEVFormer] and BEVDet4D-Occ [bevdet4d]. To ensure a fair comparison, all baselines and OccLinker are trained on the same ground truth data using identical training protocols. As shown in Table III, integrating OccLinker into existing baselines like MonoScene [MonoScene] and SurroundOcc [surroundOcc] leads to notable performance gains. In particular, the combination of OccLinker with SurroundOcc achieves the best results among all models, improving IoU and mIoU (All) by 1.63% and 0.37%, respectively. Moreover, compared with BEVFormer, which incorporates historical BEV features, our method further boosts performance by 0.41% in IoU and 0.34% in mIoU.
Occupancy accuracy on Occ3D benchmark
We also evaluate our method on the Occ3D benchmark [Occ3D], as shown in Table IV. The baselines include single-frame VON methods such as MonoScene [MonoScene], OccFormer [OccFormer], SurroundOcc [surroundOcc], and ViewFormer [viewformer], as well as history-aware VON methods including SparseOcc [SparseOcc_Liu], BEVDet4D-Occ [bevdet4d], and OPUS-L [opus]. When integrated into single-frame VONs, OccLinker consistently improves the performance of the base models. Furthermore, in ViewFormer+OccLinker, ViewFormer maintains a memory queue in which each frame’s intermediate BEV feature interacts with historical multi-frame BEV representations, while our method further supplements the missing sparse motion features between frames. As a result, ViewFormer+OccLinker surpasses ViewFormer by 0.24% in IoU and 0.94% in mIoU (All).
We note that in Table IV, ViewFormer achieves higher IoU than SurroundOcc, while this advantage is not reflected in $\overline{S_{s}}$. This is mainly because ViewFormer attains higher IoU by aggregating multi-frame BEV features, which captures richer geometric context for individual frames. However, its complex temporal fusion may occasionally propagate redundant or noisy historical information, resulting in less stable transitions between consecutive frames. In contrast, SurroundOcc performs frame-by-frame prediction with limited reliance on long-term history, which naturally leads to smoother temporal behavior and thus higher temporal consistency scores, albeit at the cost of lower per-frame spatial accuracy. Notably, our method improves the temporal consistency of both categories of models by explicitly enhancing sparse inter-frame correlations without introducing additional temporal noise.
Method IoU $\uparrow$ mIoU $\uparrow$ $\overline{S_{\rm m}}\uparrow$ $\overline{S_{\rm s}}\uparrow$
@@
Single-frame VON methods
Atlas [Atlas] 28.66 15.00 12.64 17.35
TPVFormer [TPVFormer] 30.86 17.10 14.04 20.15
MonoScene [MonoScene] 10.04 1.15 0.24 2.07 46.53 81.77
+OccLinker $\Delta$ 13.10 +3.06 1.69 +0.54 0.34 +0.10 3.04 +0.97 54.21 +7.68 83.84 +2.07
SurroundOcc [surroundOcc] 31.49 20.30 18.39 22.20 58.33 91.71
+OccLinker $\Delta$ 33.12 +1.63 20.67 +0.37 18.26 -0.13 23.08 +0.88 60.64 +2.31 92.54 +0.83
History-aware VON methods
BEVDet4D-Occ [bevdet4d] 24.26 14.22 11.10 17.34
BEVFormer [BEVFormer] 30.50 16.75 14.17 19.33 52.84 84.21
+OccLinker $\Delta$ 30.91 +0.41 17.09 +0.34 14.16 -0.01 20.01 +0.68 55.21 +2.37 85.99 +1.78
Table III: Occupancy prediction performance on the SurroundOcc benchmark [surroundOcc]. Performance gains ($\Delta$) are highlighted in bold, and the best results are marked with .
Method IoU $\uparrow$ mIoU $\uparrow$ $\overline{S_{\rm m}}\uparrow$ $\overline{S_{\rm s}}\uparrow$
@@
Single-frame VON methods
MonoScene [MonoScene] 6.06 5.36 6.68
OccFormer [OccFormer] 21.93 21.78 22.06
SurroundOcc [surroundOcc] 51.89 7.24 0.36 13.35 65.35 89.54
+OccLinker $\Delta$ 52.13 +0.24 10.33 +3.09 1.98 +1.62 17.76 +4.41 69.60 +4.25 90.91 +1.37
History-aware VON methods
SparseOcc [SparseOcc_Liu] 30.10
BEVDet4D-Occ [bevdet4d] 39.30 29.09 42.16
OPUS-L [opus] 36.20 31.25 40.44
ViewFormer [viewformer] 70.39 40.46 33.73 46.45 67.26 86.06
+OccLinker $\Delta$ 70.63 +0.24 41.30 +0.84 34.33 +0.60 47.50 +1.05 70.13 +2.87 87.10 +1.04
Table IV: Occupancy prediction performance on the Occ3D benchmark [surroundOcc]. Performance gains ($\Delta$) are highlighted in bold, and the best results are marked with .
@@
The $\overline{S_{\rm m}}$ and $\overline{S_{\rm s}}$ results presented in Tables III and IV show that integrating OccLinker consistently improves the temporal consistency of occupancy predictions across all frames and scenes for all models, highlighting the effectiveness of our approach. This improvement can be attributed to the use of previous static features and the incorporation of motion information, which together provide valuable correlations between static/dynamic cues and occupancy refinement.
@@ IV-D Ablation study
Table V presents the ablation results of different decomposed correlation designs in OccLinker, as illustrated in Fig. 4, when integrated with SurroundOcc [surroundOcc] under $L=1$. M0 denotes the baseline model without OccLinker. M1 removes the correlation between historical and current static features. M2 omits the current static features entirely. M3 excludes the correlation between motion features and current static features. M4 reverses the roles of queries and key/value pairs in the correlation design of Fig. 4. M4 represents the complete version of OccLinker with all components included.
@@
Table VI investigates the impact of different types of motion information on SurroundOcc+OccLinker when $L=1$. Specifically, I0 denotes the base model [surroundOcc] without any motion input. I1 uses raw high-frequency intermediate frames as motion input for OccLinker. I2 incorporates optical flow [OpticalFlow], while I3 employs frame difference [Frame_difference] to extract motion cues.
Idx. Motion type IoU $\uparrow$ mIoU $\uparrow$ $\overline{S_{\rm m}}\uparrow$ $\overline{S_{\rm s}}\uparrow$ Time $\downarrow$
@@
As shown in Table VI, I1 improves IoU over I0, but performs worst in mIoU, $\overline{S_{\rm m}}$, and $\overline{S_{\rm s}}$ among the motion-aware variants. This performance drop is largely due to the presence of redundant and irrelevant information in raw intermediate frames, which complicates the extraction of motion features by OccLinker. Moreover, I3 significantly outperforms I2 in both IoU and mIoU, and achieves slightly better results in $\overline{S_{\rm m}}$ and $\overline{S_{\rm s}}$. This suggests that frame difference is more effective at capturing abrupt scene changes, such as pedestrians suddenly appearing or vehicles exiting intersections, whereas optical flow may introduce delays in responding to such dynamics. In addition, thanks to the lightweight design of OccLinker, using frame difference further reduces computational overhead by operating on simple differential signals, thereby enhancing inference efficiency.
@@
We conduct ablation experiments on the temporal context window length $L\in\{0,1,2,3,4\}$ in $\mathcal{W}$. The case of $L=0$ corresponds to SurroundOcc [surroundOcc], which does not utilize any historical keyframes. As shown in Table VII, the best IoU and mIoU are achieved when $L=1$, while the highest $\overline{S_{\rm m}}$ and $\overline{S_{\rm s}}$ scores occur at $L=2$.
Idx. Window length $L$ IoU $\uparrow$ mIoU $\uparrow$ $\overline{S_{\rm m}}\uparrow$ $\overline{S_{\rm s}}\uparrow$
@@
When introducing historical keyframes (from $L=0$ to $L=1$), both occupancy accuracy metrics (IoU and mIoU) and temporal consistency metrics ($\overline{S_{\rm m}}$ and $\overline{S_{\rm s}}$) are improved, indicating that OccLinker enables SurroundOcc to leverage additional historical information and thus enrich temporal context. However, when $L$ is further increased to 2, although temporal consistency continues to improve, larger values of $L$ also introduce more irrelevant or noisy features. This makes it more challenging to associate historical information with the current occupancy prediction and may potentially lead to semantic misalignment.
@@ IV-E Case analysis
To qualitatively assess the effectiveness of our method, we compare SurroundOcc+OccLinker with the a single-frame 3D VON baseline (SurroundOcc [surroundOcc]) and a history-aware VON method (BEVDet4D-Occ [bevdet4d]).
@@
As shown in Fig. 5 (Scene 277, Frames #7–#11), a pedestrian walking along the sidewalk – parallel to the ego vehicle’s trajectory – is intermittently occluded by roadside vegetation. SurroundOcc [surroundOcc] exhibits significant instability, with the pedestrian missing in Frames #7 and #9, highlighting its limitations in temporal modeling. BEVDet4D-Occ [bevdet4d] partially alleviates this issue by fusing historical latent features into the current latent space, but still shows occasional inconsistencies, such as a detection dropout in Frame #8. In contrast, SurroundOcc+OccLinker completely eliminates these flickering artifacts and achieves consistent detection across all occlusion states.
@@
Fig. 7 showcases several challenging scenarios: (i) Vehicle-pedestrian occlusion (Scene-0911 Frame #15, Scene-0928 Frame #14): Both SurroundOcc [surroundOcc] and BEVDet4D-Occ [bevdet4d] fail to recover the occupancy state of the partially occluded pedestrian, whereas our method accurately localizes the target with precise geometry. (ii) Ambiguous road boundary (Scene-0923 Frame #28): Baseline methods produce fragmented or erroneous occupancy predictions. In contrast, our approach correctly anticipates the right-turn road geometry and preserves superior shape consistency with the actual road layout.
@@ IV-F Efficiency–performance tradeoff analysis
To ensure a fair comparison, all overhead analysis experiments are conducted on a single NVIDIA L20 GPU. Table Table IX summarizes the results of several state-of-the-art methods, including history-aware approaches (FB-Occ [fb_occ], SparseOcc [SparseOcc_Liu], OPUS-L/M/S/T [opus], and BEVDet4D-Occ [bevdet4d]), the single-frame 3D VON baseline (ViewFormer [viewformer]), and our method combined with ViewFormer. All models are evaluated on the Occ3D benchmark [Occ3D] using the same 2D visual encoder backbone, ResNet-50, which outputs feature maps with a resolution of $(32,88)$. The table reports the mIoU, GPU memory usage during inference, and per-sample latency of each method.
@@
FB-Occ [fb_occ] 39.11 5,933 MB 0.09 s
SparseOcc [SparseOcc_Liu] 30.10 7,147 MB 0.05 s
OPUS-L [opus] 36.20 10,579 MB 0.16 s
OPUS-M [opus] 35.60 8,912 MB 0.11 s
OPUS-S [opus] 34.20 7,320 MB 0.07 s
OPUS-T [opus] 33.20 6,711 MB 0.03 s
BEVDet4D-Occ [bevdet4d] 39.30 4,689 MB 0.26 s
ViewFormer [viewformer] 40.46 4,582 MB 0.11 s
+OccLinker $\Delta$ 41.30 +0.84 4,687 MB +105 MB 0.12 s +0.01 s
@@ Acknowledgments
[Uncaptioned image] Fengcheng Yu received a bachelor’s degree in engineering from the School of Intelligent Engineering of Sun Yat-sen University, majoring in intelligent science and technology, 2025. He is currently pursuing the MS in Computer Science degree with the Viterbi School of Engineering, University of Southern California, California, USA. His research interests includes autonomous driving, high performance computing.
[Uncaptioned image] Haoran Xu received the master’s degree in Software Engineering from Guangdong University of Technology, Guangzhou, China, in 2022. He is currently pursuing the Ph.D. degree with the School of Intelligent Systems Engineering, Sun Yat-Sen University, Shenzhen, China, and also with Peng Cheng Laboratory, Shenzhen, China. His research interests include multi-agent systems, reinforcement learning, and their applications in connected and autonomous vehicles.
[Uncaptioned image] Canming Xia received the master’s degree in Control Science and Engineering from Guangdong University of Technology, Guangzhou, China, in 2023. Presently, He is pursuing the Ph.D. degree with the School of Intelligent Systems Engineering, Sun Yat-Sen University, Shenzhen, China. His research interests include decision for autonomous vehicles, reinforcement learning, and representation learning.
[Uncaptioned image] Ziyang Zong received the master’s degree in Computer Science and Technology from Nanjing University of Information Science and Technology, Nanjing, China, in 2023. Presently, she is pursuing the Ph.D. degree with the School of Intelligent Systems Engineering, Sun Yat-Sen University, Shenzhen, China. Her research interests include domain adaptation, 3D-generation, and representation learning.
[Uncaptioned image] Guang Tan (Member, IEEE) received the Ph.D. degree in computer science from the University of Warwick, Coventry, U.K., in 2007 He was a Professor with the Shenzhen Institutes of Advanced Technology, Chinese Academy of Sciences, Shenzhen, China. He has been a Professor with the School of Intelligent Systems Engineering, Sun Yat-sen University, Guangzhou, China, since 2018, where he works on networked systems and machine learning.